# LoRA-Contextualizing Adaptation of Large Multimodal Models for Long Document Understanding

### v1

@@ 5 Experiments
We assess the performance of LoCAL in evidence page retrieval and visual question answering capabilities. We first evaluate the retrieval accuracy of the Col-retrieval module within LoCAL and compare it with several baselines on SlideVQA (Tanaka et al., 2023), MM-LongBench (Ma et al., 2024c), DUDE (Van Landeghem et al., 2023), DocVQA (Mathew et al., 2020; 2021) and LoCAL-Bench. We then conduct experiments on question answering using LoCAL and compare the results with other LMM baselines, inlcuding single-page and cross-page VQA. All experiments are implemented with PyTorch and conducted on Nvidia A100 GPUs. The Col-retrieval modules are fine-tuned for 4 epochs with a batch size of 32 and a learning rate of 5e-5, using the AdamW optimizer and LoRA adapters on all linear layers in the LLM. The LoRA rank is set to 32.
@@ 5.1 Datasets
We train our Col-retrieval modules using the original training data of ColPali (Faysse et al., 2024), which includes 39,463, 10,074, 13,251, 10,000, and 45,940 question-page pairs filtered from DocVQA, InfoVQA (Mathew et al., 2022), TATDQA (Zhu et al., 2024), arXivQA (Li et al., 2024), and synthetic data across various topics, including government reports, healthcare, artificial intelligence, and energy. We also combined DocMatix-IR (Ma et al., 2024a) and PFL-DocVQA (Tito et al., 2023b) for training. We fine-tuned our QA modules using the training split of the SlideVQA dataset (Tanaka et al., 2023). The SlideVQA dataset contains 1,919 slides in the training set, 400 in the test set, and 300 in the development set, with each slide consisting of 20 pages. The training split includes 10,290 samples, each annotated with questions, answers, and corresponding evidence.
@@
We evaluated our method’s performance on four public datasets—SlideVQA, MMLongBench-Doc (Ma et al., 2024c), DocVQA (Mathew et al., 2021), and DUDE (Van Landeghem et al., 2023)—along with our proposed LoCAL-Bench dataset. The evaluation was conducted in both single-evidence (SP) and cross-evidence (MP) settings, where questions require information from either a single page or multiple pages within a long document. For DocVQA, we used 5,349 SP and 5,187 MP QA pairs from the validation split. Similarly, we combined the test and dev splits of SlideVQA to form 2,995 SP and 763 MP QA pairs for evaluation. For DUDE, we evaluated 6,307 QA pairs from the validation split.
MMLongBench-Doc, which consists of 135 PDF documents averaging 50.4 pages (ranging from 9 to 468 pages), contains 1,081 QAs in total. From these, we extracted 488 single-evidence QAs to assess the performance of large multimodal models (LMMs) designed for single-image tasks. Additionally, we report the results of our best-performing model across all categories in MMLongBench-Doc, providing a comprehensive comparison against state-of-the-art LMMs.
@@ 5.2 Evaluation Metrics
We evaluate our model’s performance on evidence retrieval and question-answering using several key metrics: Top-k Accuracy, Exact Match (EM) (Tanaka et al., 2023), Generalized Accuracy (G-Acc) (Ma et al., 2024c), Average Normalized Levenshtein Similarity (ANLS) (Biten et al., 2019), and Partial Normalized Levenshtein Similarity (PNLS) (Chen et al., 2024b). A detailed explanation of each metric can be found in Appendix B.
@@ 5.3 Comparative Retrieval Accuracy Analysis
We evaluated the accuracy of the Col-retrieval module in SlideVQA, MMLongBench-Doc, SP-DocVQA, and LoCAL-Bench, comparing it with the baseline methods including CLIP (Radford et al., 2021), BM25 (Robertson et al., 2009), and SBERT (Reimers & Gurevych, 2019). For CLIP, we used its text and image encoders to compute the cosine similarity between the feature of the question and the page. BM25 and SBERT (text-based), first used Paddle-OCR  [^3] PaddleOCR: https://github.com/PaddlePaddle/PaddleOCR to extract page text. For SBERT, we computed feature vectors for each OCR-text box in a page and used the highest similarity score for page relevance. For BM25, the extracted text was concatenated into a single caption for relevance scoring. The results are shown in Table 1.
SlideVQA MMLong LoCAL-B SP-DocVQA
@@
Table 1: Retrieval accuracy results on four datasets, where MMLong refers to MMLongBench-Doc, LoCAL-B refers to LoCAL-Bench. Bold font indicates the best model.
@@ 5.4 Main Results
Table 2 presents the comparison results. We first evaluate LoCAL on single-evidence questions from SP-SlideVQA, MMLongBench-Doc, and SP-DocVQA, where the required information is on a single page. To demonstrate the question-answering capabilities of LMMs, we include four “cheating” baselines where models are given the ground truth evidence page. Next, we test LoCAL on cross-evidence questions from MP-SlideVQA, MP-DocVQA, and DUDE, where information spans multiple pages. We only test LoCAL with InternVL2-4B backbone, since the other two LMM are pre-trained for single-page understanding. LoCAL’s performance is compared with classical encoder-only and encoder-decoder models, including BERT (Kenton & Toutanova, 2019), Longformer (Beltagy et al., 2020), Big Bird (Zaheer et al., 2020), T5 (Raffel et al., 2020), Hi-VT5 (Tito et al., 2023a), and LayoutLMv3 (Huang et al., 2022), with results taken from the best settings in the original papers. InternVL2-8B and GPT-4o, processing all pages, serve as the state-of-the-art baselines for open-source and proprietary multipage LMMs, respectively. Finally, we demonstrate how the limitations of the retrieval and QA modules can impact overall performance through challenging examples from the SlideVQA dataset, as shown in Appendix C.
@@
Multi-image LMMs
@@
LoCAL Models (Proposed)
LoCAL-PaliGemma 3B R1 35.03 0.60 23.9 0.35 0.56 0.69
LoCAL-PaliGemma 3B R1 49.75 0.65 23.1 0.38 0.56 0.68
LoCAL-Phi-3-V 4B R1 12.85 0.78 30.7 0.50 0.55 0.75
LoCAL-Phi-3-V 4B R1 58.13 0.77 28.4 0.44 0.68 0.73
LoCAL-InternVL2 4B R5 16.40 0.58 33.2 0.48 0.70 0.76
LoCAL-InternVL2 4B R5 45.07 0.77 34.0 0.49 0.71 0.75
@@
Multi-image LMMs
@@
LoCAL Models (Proposed)
LoCAL-InternVL2 4B R5 24.25 0.61 0.70 0.76 0.36 0.57
LoCAL-InternVL2 4B R5 31.98 0.59 0.71 0.76 0.45 0.54
@@
We observe LoCAL consistently outperforms InternVL2-8B, across various settings. The primary issue with LMMs is that long documents are transformed into excessively long visual token sequences, leading to significant memory burdens, as reported later in section 5.5. In datasets like MMLongBench-Doc and DocVQA, some documents exceed hundreds of pages, causing out-of-memory errors, even on servers with 8 $\times$ A100 (80GB) GPUs. In such cases, we assigned a zero score in our experiments. In contrast, GPT-4o exhibits strong multi-page reasoning capabilities. However, the accuracy of the cheating baseline slightly surpasses that of using all pages, as providing only the evidence pages helps GPT-4o avoid distractions from irrelevant information in the longer context. Moreover, LoCAL with InternVL2-4B backbone perform slightly better than the one with Phi-3-V backbone on MMLongBench-Doc and SP-DocVQA, possibly due to the improvement in retrieval accuracy by using top-5 pages, which is more crucial for longer documents.
@@
We observe that LoCAL QA modules with PaliGemma and InternVL2-4B backbones show a significant increase in EM on the SlideVQA dataset, surpassing their cheating baselines after fine-tuning on SlideVQA. The model with the Phi-3-V backbone shows notable improvements in Exact Match (EM) scores without gains in PNLS, suggesting that fine-tuning primarily enhanced the model’s attention and answer formatting. This could be because the pre-trained model was already optimized for these question types. Nevertheless, as shown in Figure D.1, we empirically find that fine-tuning still improves answering performance. However, we notice a performance drop for fine-tuned LoCAL-Phi-3-V on MMLongBench-Doc, indicating that fine-tuning can harm LLM generalization. A similar trend is seen with the InternVL2-4B backbone on the DUDE dataset.
@@
Finally, we present the complete results of LoCAL-InternVL2-4B on the MMLongBench-Doc dataset to highlight the advantages of our method. As shown in Table 3, our model, with only 4 billion parameters, outperforms all open-source LMMs and achieves performance comparable to proprietary models such as Claude-3 Opus and Gemini-1.5-Pro.
@@
LoCAL Models (Proposed)
LoCAL-InternVL2 (R5) 4B 26.5 18.8 22.3 19.6 23.6 33.2 13.1 12.4 22.2 22.8
LoCAL-InternVL2 (R5) 4B 26.3 22.1 25.0 20.7 25.2 34.0 10.6 15.7 23.0 24.2
@@
Table 3: Performance of various models on MMLongBench-Doc. Questions are categorized in two ways: (1) by evidence source type—text (TXT), layout (LAY), chart (CHA), table (TAB), and image (IMG); and (2) by evidence pages—single-page (SIN), cross-page (MUL), and unanswerable (UNA). Models using LoRA adapters fine-tuned on SlideVQA for the QA module are marked with $\dagger$. Bold font indicates the best open-source model. The results of baseline models are adopted from the original MMLongBench-Doc paper Ma et al. (2024c).
@@ 5.5 Efficiency of Different Models
To evaluate the efficiency of LoCAL, we conducted experiments on the SlideVQA dataset, which has 20 pages per question with a resolution of 1024x768. We recorded peak GPU memory usage and time costs for retrieval and QA modules separately. The GPU memory is manually recorded using the nvidia-smi command, which tends to report higher numbers than the actual memory required by the application due to overhead and memory management processes. We tested backbones including PaliGemma, Phi-3-v, and InternVL2-4B, all equipped with LoRA adapters. Since PaliGemma and Phi-3-v are single-page models, we used top-1 retrieved image as input. InternVL2-4B, however, supports multi-image input, allowing us to test with the top-1, 5, and 12 retrieved images.
LoCAL-Backbone Page Retrieval QA
@@
Table 4: Time (s) cost and Peak GPU memory (GB) cost of LoCAL models with different backbones.
As shown in Table 4, the memory consumption of the QA module increases with the number of evidence pages, with 13 images (1024x768) exceeding the 80GB limit on an A100 GPU resulting in out-of-memory error. In contrast, the memory usage of the retrieval module remains low, as LoCAL processes pages independently during page retrieval, with memory costs equivalent to single-page reasoning. Despite the higher memory demands for multi-evidence QA, LoCAL remains compact and time-efficient, making it well-suited for answering questions from fewer evidence pages in resource-constrained environments. This demonstrates LoCAL’s ability to balance performance and resource usage, ensuring scalability across diverse deployment scenarios.
@@ 6 Conclusions
In this paper, we propose LoCAL, a lightweight LMMs for visually-rich document understanding. LoCAL has a unique design to facilitate multi-page document understanding using dual LoRA adapters. The research highlights that small open-source models are great at processing multipage documents and underscored the importance of efficient retrieval mechanisms in filtering irrelevant pages. Furthermore, we collect the LoCAL-bench dataset for document understanding, and empirical results on benchmarks demonstrated the effectiveness of LoCAL. We hope these findings provide valuable insights for optimizing lightweight LMMs, aiming to improve accuracy and efficiency in visually-rich document understanding.
@@ 7 Limitations
LoCAL is the first LMM that can perform retrieval and question answering simultaneously. However, it still requires computational resources for training and inference, which may limit its practical applicability in resource-constrained environments. LoCAL should be mobile friendly, as it only requires a single base model. This base model can be Phi-3-Silica within MS operating systems or an Apple on-device model within IOS 18. A routing mechanism in Apple Intelligence can better balance computational cost and performance. However, our experiments are not performed on these real-world cases, which are useful for pushing forward document intelligence.

### v2

@@ 5 Experiments
We assess the performance of SV-RAG in evidence page retrieval and visual question answering capabilities. We first evaluate the retrieval accuracy of the Col-retrieval module within SV-RAG and compare it with several baselines on SlideVQA (Tanaka et al., 2023), MM-LongBench (Ma et al., 2024d), DUDE (Van Landeghem et al., 2023), DocVQA (Mathew et al., 2020; 2021) and VisR-Bench. We then conduct experiments on question answering using SV-RAG and compare the results with other LMM baselines, inlcuding single-page and cross-page VQA. All experiments are implemented with PyTorch and conducted on Nvidia A100 GPUs. The Col-retrieval modules are fine-tuned for 4 epochs with a batch size of 32 and a learning rate of 5e-5, using the AdamW optimizer and LoRA adapters on all linear layers in the LLM. The LoRA rank is set to 32.
@@ 5.1 Datasets
We train our Col-retrieval modules using the original training data of ColPali (Faysse et al., 2024), which includes 39,463, 10,074, 13,251, 10,000, and 45,940 question-page pairs filtered from DocVQA, InfoVQA (Mathew et al., 2022), TATDQA (Zhu et al., 2024), arXivQA (Li et al., 2024), and synthetic data across various topics, including government reports, healthcare, artificial intelligence, and energy. We incorporated DocMatix-IR (Ma et al., 2024b) and PFL-DocVQA (Tito et al., 2023b), using GPT-4o to filter out duplicate images and unsuitable questions. The expanded dataset improved top-1 retrieval accuracy on MMLongBench-Doc by $\sim$1without affecting other benchmarks. We fine-tuned our QA modules using the training split of the SlideVQA dataset (Tanaka et al., 2023). The SlideVQA dataset contains 1,919 slides in the training set, 400 in the test set, and 300 in the development set, with each slide consisting of 20 pages. The training split includes 10,290 samples, each annotated with questions, answers, and corresponding evidence.
@@
We evaluated our method’s performance on four public datasets—SlideVQA, MMLongBench-Doc (Ma et al., 2024d), DocVQA (Mathew et al., 2021), and DUDE (Van Landeghem et al., 2023)—along with our proposed VisR-Bench dataset. The evaluation was conducted in both single-evidence (SP) and cross-evidence (MP) settings, where questions require information from either a single page or multiple pages within a long document. For DocVQA, we used 5,349 SP and 5,187 MP QA pairs from the validation split. Similarly, we combined the test and dev splits of SlideVQA to form 2,995 SP and 763 MP QA pairs for evaluation. For DUDE, we evaluated 6,307 QA pairs from the validation split.
MMLongBench-Doc, which consists of 135 PDF documents averaging 50.4 pages (ranging from 9 to 468 pages), contains 1,081 QAs in total. From these, we extracted 488 single-evidence QAs to assess the performance of MLLMs designed for single-image tasks. Additionally, we report the results of our best-performing model across all categories in MMLongBench-Doc, providing a comprehensive comparison against state-of-the-art LMMs.
@@ 5.2 Evaluation Metrics
We evaluate our model’s performance on evidence retrieval and question-answering using several key metrics: Top-k Accuracy, Exact Match (EM) (Tanaka et al., 2023), Generalized Accuracy (G-Acc) (Ma et al., 2024d), Average Normalized Levenshtein Similarity (ANLS) (Biten et al., 2019), and Partial Normalized Levenshtein Similarity (PNLS) (Chen et al., 2024b). A detailed explanation of each metric can be found in Appendix B.
@@ 5.3 Comparative Retrieval Accuracy Analysis
We evaluated the accuracy of the Col-retrieval module in SlideVQA, MMLongBench-Doc, SP-DocVQA, and VisR-Bench, comparing it with the baseline methods including CLIP (ViT-L/14) (Radford et al., 2021), SigLip (so400m-patch14-384) Zhai et al. (2023), BM25 (Robertson et al., 2009), SBERT (Reimers & Gurevych, 2019), BGE-M3 Chen et al. (2024c), BGE-large Xiao et al. (2023), and NV-Embed-v2 Lee et al. (2024). For encoder models, we used their text and image encoders to compute the cosine similarity between the feature of the question and the page. For text-based methods, the text content in the MMLongBench-Doc and SV-RAG Bench datasets is extracted using a document parser to ensure higher accuracy. For SlideVQA and SP-DocVQA, where only scanned images are available, the text is extracted using Paddle-OCR [^2] PaddleOCR: https://github.com/PaddlePaddle/PaddleOCR.
SlideVQA MMLong VisR-B SP-DocVQA
@@
BGE-M3 74.3 92.0 42.7 66.6 47.7 78.1 47.8 77.5
Bge-large 81.3 93.3 47.4 71.5 53.7 80.3 56.7 81.5
NV-Embed-v2 82.2 94.3 47.4 69.0 55.2 82.7 51.7 80.2
Encoder Models
CLIP 58.4 86.9 32.4 63.4 33.4 62.1 37.1 69.4
SigLip 66.2 90.1 44.9 69.4 53.2 81.3 39.3 71.9
@@
Col-Phi-3-vision 90.6 98.8 64.8 84.8 71.9 91.8 65.1 87.0
Table 1: Retrieval accuracy results on four datasets, where MMLong refers to MMLongBench-Doc, SV-RAG-B refers to SV-RAG-Bench. Bold font indicates the best model.
@@ 5.4 Main Results
Table 2 presents the comparison results. We first evaluate SV-RAG on single-evidence questions from SP-SlideVQA, MMLongBench-Doc, and SP-DocVQA, where the required information is on a single page. To demonstrate the question-answering capabilities of LMMs, we include four “cheating” baselines where models are given the ground truth evidence page. Next, we test SV-RAG on cross-evidence questions from MP-SlideVQA, MP-DocVQA, and DUDE, where information spans multiple pages. We only test SV-RAG with InternVL2-4B backbone, since the other two LMM are pre-trained for single-page understanding. SV-RAG’s performance is compared with classical encoder-only and encoder-decoder models, including BERT (Kenton & Toutanova, 2019), Longformer (Beltagy et al., 2020), Big Bird (Zaheer et al., 2020), T5 (Raffel et al., 2020), Hi-VT5 (Tito et al., 2023a), and LayoutLMv3 (Huang et al., 2022), with results taken from the best settings in the original papers. InternVL2-8B and GPT-4o, processing all pages, serve as the state-of-the-art baselines for open-source and proprietary multipage LMMs, respectively. We demonstrate how the limitations of the retrieval and QA modules can impact overall performance through challenging examples from the SlideVQA dataset, as shown in Appendix C. Additional comparisons with text-only baselines utilizing a document parser are provided in Appendix G.1.
@@
Multi-image MLLMs
@@
SV-RAG Models (Proposed)
SV-RAG-PaliGemma 3B R1 35.03 0.60 23.9 0.35 0.56 0.69
SV-RAG-PaliGemma 3B R1 49.75 0.65 23.1 0.38 0.56 0.68
SV-RAG-Phi-3-vision 4B R1 12.85 0.78 30.7 0.50 0.55 0.75
SV-RAG-Phi-3-vision 4B R1 58.13 0.77 28.4 0.44 0.68 0.73
SV-RAG-InternVL2 4B R5 16.40 0.58 33.2 0.48 0.70 0.76
SV-RAG-InternVL2 4B R5 45.07 0.77 34.0 0.49 0.71 0.75
@@
Multi-image MLLMs
@@
SV-RAG Models (Proposed)
SV-RAG-InternVL2 4B R5 24.25 0.61 0.70 0.76 0.36 0.57
SV-RAG-InternVL2 4B R5 31.98 0.59 0.71 0.76 0.45 0.54
@@
We observe SV-RAG consistently outperforms InternVL2-8B, across various settings. The primary issue with LMMs is that long documents are transformed into excessively long visual token sequences, leading to significant memory burdens, as reported later in section 5.5. In datasets like MMLongBench-Doc and DocVQA, some documents exceed hundreds of pages, causing out-of-memory errors, even on servers with 8 $\times$ A100 (80GB) GPUs. In such cases, we assigned a zero score in our experiments. In contrast, GPT-4o exhibits strong multi-page reasoning capabilities. However, the accuracy of the cheating baseline slightly surpasses that of using all pages, as providing only the evidence pages helps GPT-4o avoid distractions from irrelevant information in the longer context. Moreover, SV-RAG with InternVL2-4B backbone perform slightly better than the one with Phi-3-vision backbone on MMLongBench-Doc and SP-DocVQA, possibly due to the improvement in retrieval accuracy by using top-5 pages, which is more crucial for longer documents.
@@
We observe that SV-RAG QA modules with PaliGemma and InternVL2-4B backbones show a significant increase in EM on the SlideVQA dataset, surpassing their cheating baselines after fine-tuning on SlideVQA. The model with the Phi-3-vision backbone shows notable improvements in Exact Match (EM) scores without gains in PNLS, suggesting that fine-tuning primarily enhanced the model’s attention and answer formatting. This could be because the pre-trained model was already optimized for these question types. Nevertheless, as shown in Figure D.1, we empirically find that fine-tuning still improves answering performance. However, we notice a performance drop for fine-tuned SV-RAG-Phi-3-vision on MMLongBench-Doc, indicating that fine-tuning can harm LLM generalization. A similar trend is seen with the InternVL2-4B backbone on the DUDE dataset.
@@
Finally, we present the complete results of SV-RAG-InternVL2-4B on the MMLongBench-Doc dataset to highlight the advantages of our method. As shown in Table 3, our model, with only 4 billion parameters, outperforms all open-source LMMs and achieves performance comparable to proprietary models such as Claude-3 Opus and Gemini-1.5-Pro.
@@
SV-RAG Models (Proposed)
SV-RAG-InternVL2 (R5) 4B 26.5 18.8 22.3 19.6 23.6 33.2 13.1 12.4 22.2 22.8
SV-RAG-InternVL2 (R5) 4B 26.3 22.1 25.0 20.7 25.2 34.0 10.6 15.7 23.0 24.2
@@
Table 3: Performance of various models on MMLongBench-Doc. Questions are categorized in two ways: (1) by evidence source type—text (TXT), layout (LAY), chart (CHA), table (TAB), and image (IMG); and (2) by evidence pages—single-page (SIN), cross-page (MUL), and unanswerable (UNA). Models using LoRA adapters fine-tuned on SlideVQA for the QA module are marked with $\dagger$. Bold font indicates the best open-source model. The results of baseline models are adopted from the original MMLongBench-Doc paper Ma et al. (2024d).
@@ 5.5 Efficiency of Different Models
To evaluate the efficiency of SV-RAG, we conducted experiments on the SlideVQA dataset, which has 20 pages per question with a resolution of 1024x768. We recorded peak GPU memory usage and time costs for retrieval and QA modules separately. The GPU memory is manually recorded using the nvidia-smi command, which tends to report higher numbers than the actual memory required by the application due to overhead and memory management processes. We tested backbones including PaliGemma, Phi-3-v, and InternVL2-4B, all equipped with LoRA adapters. Since PaliGemma and Phi-3-v are single-page models, we used top-1 retrieved image as input. InternVL2-4B, however, supports multi-image input, allowing us to test with the top-1, 5, and 12 retrieved images.
SV-RAG-Backbone Page Retrieval QA
@@
Table 4: Time (s) cost and Peak GPU memory (GB) cost of SV-RAG models with different backbones.
As shown in Table 4, the QA module’s memory consumption increases with the number of evidence pages, with 13 images (1024x768) exceeding the 80GB limit on an A100 GPU resulting in out-of-memory error. In contrast, the retrieval module maintains low memory usage, as SV-RAG processes pages independently, with costs equivalent to single-page reasoning. Although multi-evidence QA requires more memory, SV-RAG remains efficient and compact, making it well-suited for answering questions from fewer evidence pages in resource-constrained environments. This demonstrates SV-RAG’s ability to balance performance and resource usage, ensuring scalability across diverse deployment scenarios. Additional results on the retrieval efficiency are presented in Table F.1.
@@ 6 Conclusions
In this paper, we propose SV-RAG, a lightweight MLLMs for visually-rich document understanding. SV-RAG has a unique design to facilitate multi-page document understanding using dual LoRA adapters. The research highlights that small open-source models are great at processing multipage documents and underscored the importance of efficient retrieval mechanisms in filtering irrelevant pages. Furthermore, we collect the VisR-bench dataset for document understanding, and empirical results on benchmarks demonstrated the effectiveness of SV-RAG. We hope these findings provide valuable insights for optimizing lightweight MLLMs, aiming to improve accuracy and efficiency in visually-rich document understanding.
@@ 7 Limitations
SV-RAG is the first MLLM that can perform visual retrieval-augmented generation for document question answering using a single model. However, it still requires computational resources for training and inference, which may limit its practical applicability in resource-constrained environments. SV-RAG should be mobile friendly, as it only requires a single base model. This base model can be Phi-3-Silica within MS operating systems or an Apple on-device model within Apple IOS 18. A routing mechanism in Apple Intelligence can better balance computational cost and performance. However, our experiments are not performed on these real-world devices, which are necessary for pushing forward document intelligence.
8 Ethics Statement
The VisR-Bench dataset was curated with careful consideration of ethical and legal concerns. All documents are sourced from publicly available data with licenses explicitly permitting research use. To ensure data integrity and compliance, we provide links to the original sources instead of distributing the documents. Additionally, all QA pairs have been manually reviewed to exclude harmful content and personally identifiable information. The dataset does not expose sensitive user data, and experimental results are reported as aggregate statistics to prevent information leakage while ensuring reproducibility. These measures uphold ethical and legal standards while supporting responsible AI research.